On her way to work from home, Jayla averaged only 20 mph. On her drive home, she averaged 40 mph. If the total travel time was 1.5 hours, how long did it take her to drive to work?
Answer by galileo(7)

On her way to work from home, Jayla averaged only 20 mph. On her drive home, she averaged 40 mph. If the total travel time was 1.5 hours, how long did it take her to drive to work?
Let x = time to drive to work
1.5-x = time to drive home
from your lesson in physics, distance traveled = average speed times time
d = vt
since the distance traveled in driving to work and back home are equal, we have
20(x) = 40(1.5-x)
20x = 60 - 40x
20x + 40x = 60
60x = 60
x = 60/60
x = 1 hr (time to drive to work)